OpenStudy (anonymous):

Explain how to solve a quadratic equation using the quadratic formula. Demonstrate the process with your own example.

OpenStudy (anonymous):

\[x^2+4x+3\]when plugging in numbers into the formula, u use the coefficients to each number. with a b and c, it goes in order: a= the number in front of the x^2, b= the number infront of the x and c= the number by itself. be sure to include the sign infront of each number. ex: -2x^2, a= -2. Now using \[(-b(+-)\sqrt{b^2-4ac})/2a\] a=1, b=4 c=3 so -4 plus or minus the square root of 16-12 all over 2 equals -4plus or minus 2 over 2, therefore -4-2 over 2 =-3 and -4+2=-2 over two= -1 so x=-1,-3 or u can write it (x+3)(x+1)